OpenStudy (anonymous):

at a factory that produces pistons for cars, machine 1 produced 700 satisfactory pistons and 300 unsatisfactory pistons today. machine 2 produced 707 satisfactory pistons and 303 unsatisfactory pistons today. suppose that one piston from machine 1 and one piston from machine 2 are chosen at random from today's batch. what is the probability that the piston chosen from machine 1 is unsatisfactory and the piston chosen from machine 2 is satisfactory?

OpenStudy (anonymous):

the proccess of choosing is independent so probability of choosing from machin1 is:300/1000 =0.3 choosing from machin2 is:707/1010=0.7 the total result is 0.3*0.7=0.21 ;

OpenStudy (anonymous):

thanks soo much@!

OpenStudy (anonymous):

you're welcome!